OpenStudy (anonymous):

A takes 3 times long as B+C together and B takes 4 times as long as A+C together. All 3 working together takes 24 days to complete, then time taken by A alone to complete???

OpenStudy (kc_kennylau):

Let the work A B C can complete in 1 hour be a b c respectively, set up 3 equations :)

OpenStudy (anonymous):

how can i set up equation for other two..can u show me.

OpenStudy (kc_kennylau):

1/a=3/(b+c)

OpenStudy (anonymous):

that's the time given...3 times not work done by each, right?

OpenStudy (anonymous):

@kc_kennylau

OpenStudy (perl):

still need help?

OpenStudy (kc_kennylau):

well, you can treat it as 3a=b+c, that means the work b and c can do together is three times the work a does.

OpenStudy (kc_kennylau):

\[\left\{\begin{array}{rcl} 3a&=&b+c\\ 4b&=&a+c\\ a+b+c&=&\frac1{24} \end{array}\right.\] (It is in terms of days instead of hours)

OpenStudy (dan815):

A=a work/days B = b work/days C= c work/days time taken by a to do his work w' = w/A = 3*(w/(b+c)) w/B=4*(w/(A+C)) w/(A+B+C)=24

OpenStudy (dan815):

if i am cutting the grass at 2 sqft/min, you are cutting the grass at 4sqft/min together we are cutting 6 sqftt/min so we shud be done 2*3=6, 3 times faster than me cutting alone, in other words our time together is 1/3rd the time of mine alone

OpenStudy (dan815):

same principle

OpenStudy (anonymous):

thanks!!!!